OpenStudy (anonymous):

Find the polynomial function of the lowest degree with rational coefficients that has the given numbers as some of its zeros. 4-i,square root of 12

OpenStudy (anonymous):

two other roots are \(4+i\) and \(-\sqrt{12}\)

OpenStudy (anonymous):

Irrationals and conjugates occurs in pairs.

OpenStudy (anonymous):

so your job is to multiply out \[(x+\sqrt{12})(x-\sqrt{12})(x-(4+i))(x-(4-i))\] whih is not nearly as hard as it looks